OpenStudy (anonymous):

3x + 8y = 6 -4x -9y = -2 use substitution

OpenStudy (anonymous):

can someone help ?

OpenStudy (anonymous):

does anyone help on here or is it just a chat session ?

OpenStudy (anonymous):

Substitution means that you solve for one variable and then stick the resultant value back in for the other. For example, we can the first equation for x \[3x = 6 - 8y\]\[x = \frac{6 - 8y}{3}\] Then put that in to the second one. \[-4(\frac{6 - 8y}{3}) - 9y = -2\] Once you've got y, you can find x, of course.

OpenStudy (anonymous):

solve one of the lines for a single variable and then substitute that variable in the equation you haven't already solved for

OpenStudy (anonymous):

can you explain a little different , my instructor showed me a different way but cant find the paper that I wrote it down on

OpenStudy (jusaquikie):

solve one equation for x or y, let's use x, then substitute the equation you had for the answere for x in place of the x in the other equation and solve for y, once you have y as a number plug that in to either equation and solve for x

OpenStudy (jusaquikie):

SACAPUNTAS put it pretty clearly.

OpenStudy (anonymous):

Different how? It's really just as simple as solving one equation for one variable and putting what you find back into the other equation. You could also solve for y, of course. \[y = \frac{6 - 3x}{8}\] Unless you mean a different method, but I thought we were talking about substitution. :D

OpenStudy (anonymous):

yes I was talking about substitution

OpenStudy (jusaquikie):

no, you are correct this is substitution method

OpenStudy (anonymous):

Ok. You know how to solve for a variable, right?

OpenStudy (anonymous):

thanks for the help I will figure it out on my own